OpenStudy (anonymous):

Write the equation of a line in point slope form that goes through (2 , -4) and ( -1 , 0)

OpenStudy (eamier):

you know this formula y=mx+c

OpenStudy (anonymous):

thats y intercept not point slope

OpenStudy (anonymous):

slope*

OpenStudy (eamier):

sorry forgot that

OpenStudy (anonymous):

haha no worries

OpenStudy (eamier):

is it like this \[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

yes

OpenStudy (eamier):

its me who learning now

OpenStudy (anonymous):

hahaha

OpenStudy (eamier):

sure you know the answer right

OpenStudy (anonymous):

?

OpenStudy (eamier):

i mean you know the answer already

OpenStudy (eamier):

m=(-4-0)/(2--1)=-4/3

OpenStudy (anonymous):

yeah i know the answer now :)

OpenStudy (eamier):

this? \[y+4=\frac{-4}{3}(x-2)\]

OpenStudy (eamier):

or \[y=\frac{-4}{3}(x+1)\]